OpenStudy (anonymous):

write an equation of a line that passes through the point (-2,5) and (1,2) show work!

OpenStudy (amistre64):

i spose i can muster up a point slope form.... shudder

OpenStudy (amistre64):

(2-5) = m(1--2) -3 = m(3) m = -1 .............. y = -x -3 ?? 2 = -1(1) + b 3 = b my mistake lol y = -x +3